NASA Space Station On-Orbit Status 6 March 2017 - More CubeSats Deployed Today

©Spire Global

NASA Space Station On-Orbit Status 6 March 2017 - More CubeSats Deployed Today.

Four CubeSats were deployed this morning as the crew researched fluid shifts toward the head that may affect astronaut vision. Tools were also being collected and organized today ahead of possible maintenance spacewalks.

Four CubeSats were ejected Monday morning from outside Japan's Kibo lab module using the NanoRacks CubeSat Deployer. The LEMUR-2 satellites will help monitor global ship tracking and improve weather forecasting.

Sergey Ryzhikov from Roscosmos participated in ultrasound scans of the head and neck for the long-running Fluid Shifts study. Thomas Pesquet from the European Space Agency joined Ryzhikov for the experiment to learn how to prevent upward fluid shifts that may cause lasting eye damage.

Commander Shane Kimbrough worked inside the Quest airlock today gathering spacewalk tools. Mission planners are looking at potential spacewalks to continue upgrading the International Space Station's power systems.

On-Orbit Status Report

Microgravity Expanded Stem Cells (MESC) Microscope Operations: Over the weekend the crew completed the MESC experiment, performing two days of back-to-back operations. The crew viewed the BioCells on the microscope, then sampled and preserved the BioCells which were inserted into the Minus Eighty Degree Celsius Laboratory Freezer for ISS (MELFI) for cold stowage. The crew then deconfigured the Microgravity Science Glovebox (MSG) from the MESC experiment and reconfigured the Life Science Ancillary Hardware (LSAH) for Rodent Research activities. The goal of the MESC investigation is to determine the efficiency of using a microgravity environment to accelerate expansion (replication) of stem cells for use in terrestrial clinical trials for treatment of disease.

Fluid Shifts Baseline Imaging: With assistance of an onboard operator and ground remote guidance, the 48S subject performed the second portion of the Return-45 day (R-45) Fluid Shifts experiment. Ultrasound imaging was taken of arterial and venous measures of the head and neck, cardiac, ophthalmic and portal vein, and tissue thickness of lower and upper body. Additional measurements using the Cerebral and Cochlear Fluid Pressure (CCFP), Distortion Product Otoacoustic Emissions (DPOAE), Optical Coherence Tomography (OCT), ESA's Cardiolab (CDL) Holter Arterial Blood Pressure Device, and Intraocular Pressure (IOP) using the tonometer were taken as well. Later this week measurements will be taken in the Russian Segment using these devices in conjunction with the Chibis (Lower Body Negative Pressure) device. Fluid Shifts is a joint NASA-Russian experiment that investigates the causes for severe and lasting physical changes to astronaut's eyes. Because the headward fluid shift is a hypothesized contributor to these changes, reversing this fluid shift with a lower body negative pressure device is investigated as a possible intervention. Results from this study may help to develop preventative measures against lasting changes in vision and eye damage.

NanoRack Cubesat Deployer (NRCSD) #10 Deploy: Four Lemur and 1 TechEdSat Cubesats were successfully deployed from NRCSD today, completing the first of three planned deployments for NRCSD Mission #10. The LEMUR-2 satellites are 3U CubeSats equipped with payloads to provide two primary data products: AIS Data (Maritime Domain Awareness) and GPS-RO Data (Weather). The TechEdSat is also a 3U sized cubesat whose investigation studies a new system called the Exo-Brake which uses a spacecraft's own atmospheric drag to change its velocity and adjust its approach. This investigation tests the controls, avionics and methods that may be used on future spacecraft using the Exo-Brake system including small payloads returning to Earth or landing on Mars.

Lighting Effects Meter Reading: The crew set up and configured the Light Meter hardware and took readings in the Node 1 and US Lab. The Lab still uses the legacy General Luminaire Assembly (GLA) while the lights in Node 1 have been upgraded to the Solid State Lighting Assemblies (SSLAs). The Lighting Effects experiment aims to better quantify and qualify how lighting can effect habitability of spacecraft. The light bulbs on the ISS are being replaced with a new system designed for improved crew health and wellness. The investigation studies the impact of the change from fluorescent light bulbs to solid-state light-emitting diodes (LEDs) with adjustable intensity and color and aims to determine if the new lights can improve crew circadian rhythms, sleep, and cognitive performance. Results from this investigation also have major implications for people on Earth who use electric lights.

Extravehicular Activity (EVA) Tool Configuration: The crew began configuring tools to support the EVAs being planned for later this month.

Urine Processor Assembly (UPA): UPA was operated twice today unsuccessfully. Both runs were terminated due to high conductivity. The recommendation is to perform an Advanced Recycle Filter Tank Assembly (ARFTA) drain and fill and start a new concentration cycle. Plans are being finalized to disconnect Quick Disconnect (QD) 32 and install a CWC-I (Contingency Water Container- Iodine) to capture the purge distillate. This should return conductivity to a nominal level and prevent further high conductivity incidents. The source of the contamination in the purge distillate causing high and erratic conductivity is either the Distillation Assembly (DA) or Fluids Control and Pump Assembly (FPCA). An FCPA inspection of its liquid barrier is scheduled for March 14. This inspection will either clear or indict the FCPA as the source of the Pre-Treated Urine (PTU) in the purge distillate. A spare FCPA is on board and a spare DA is manifested on OA-7.

Mobile Servicing System (MSS) Operations: Saturday, Robotics Ground Controllers maneuvered the Space Station Remote Manipulator System (SSRMS) and Special Purpose Dexterous Manipulator (SPDM) to rotate the Enhanced Orbital Replacement Unit (ORU) Temporary Platform (EOTP) and stow the SPDM on Mobile Base System (MBS) Power Data Grapple Fixture (PDGF) 2. They then walked the SSRMS off the Node 2 PDGF onto Mobile Base System (MBS) PDGF 1 and translated the Mobile Transporter (MT) from Worksite 6 (WS6) to WS2. As they were ahead of schedule, the SSRMS was then maneuvered to unstow the SPDM and position it for the start of the removal of the Robotics Refuelling Mission (RRM) payload from EXpedite the PRocessing of Experiments to Space Station (EXPRESS) Logistics Carrier 4 (ELC4). Sunday, the RRM was removed from the ELC with SPDM Arm 2. As they were ahead of schedule, Robotics Ground Controllers proceeded to install the Stratospheric Aerosol and Gas Experiment (SAGE) Nadir Viewing Platform (NVP) onto ELC 4 with SPDM Arm 1. MSS performance was nominal for all operations.

Today's Planned Activities
All activities were completed unless otherwise noted.

HRF Generic Urine Collection
HRF Generic Sample MELFI Insertion Operations
HRF Generic Frozen Blood Collection
Fluid Shifts Ultrasound 2 HRF Rack 1 Power On
HRF Generic Refrigerated Centrifuge Configuration
ENERGY Diet Log of Breakfast
Fluid Shifts OCT Baseline Setup
Atmosphere Purification Filter Assembly АФОТ-2M, initial testing and setup in Soyuz 733.
Total Organic Carbon Analyzer (TOCA) Water Recovery System (WRS) Sample Analysis
Fluid Shifts CDL Holter Arterial BP Hardware Don - Subject
HRF Generic Refrigerated Centrifuge Spin Conclude
Fluid Shifts Ultrasound Baseline Scan
DAN Experiment Ops.
Fluid Shifts Ultrasound Baseline Scan - Operator
BEAM Intermodule Ventilation (IMV) Flow Measurement Survey
HRF Generic Frozen Blood Collection Conclude Stow
Fluid Shifts CCFP Baseline Test
Water Recovery System Water Tank Drain Termination
Extravehicular Activity (EVA) Tool Configuring
Exercise Data Downlink via OCA
Fluid Shifts OCT Baseline Exam
MRM1 Fan Screen Cleaning (Group B)
Fluid Shifts DPOAE Baseline Test
Fluid Shifts Tonometry Baseline Setup
Total Organic Carbon Analyzer (TOCA) Sample Data Record
Wastewater Storage Tank Assembly (WSTA) Fill [Aborted]
Fluid Shifts Tonometry Baseline Exam
Fluid Shifts CDL Holter Arterial BP Hardware Doff - Subject
Fluid Shifts Tonometry Baseline Stow
Fluid Shifts CCFP Baseline Stow
Battery Stowage Assembly (BSA) Operation Termination
Lighting Effects Light Meter Readings Alternate 1
Daily Electro-cardiogram registration (start)
Replace pre-treat tank [Е-К] and a hose in [АСУ] system, АСУ activation after replacement
ENERGY Diet Log of Lunch
Public Affairs Office (PAO) High Definition (HD) Config Columbus Setup
24-hour Blood Pressure Recording (start)
PAO Preparation
Public Affairs Office (PAO) Event in High Definition (HD) in Columbus
Delta file prep
Micrgravity Expanded Stem Cells MSG Teardown
NanoRacks CubeSat Deployment Photos
Fluid Shifts OCT Baseline Stow
ТГК 66P transfer operations
APEX-04 Petri Plate Photo
Metal Oxide (METOX) Regeneration Initiation
Microgravity Science Glovebox Power Down [Aborted]
Body Measures Equipment Gather
Photo TV LAB Camcorder Audio Troubleshooting
Health Maintenance System (HMS) Vision Test
СОЖ maintenance
HRF Generic Sample MELFI Retrieval Insertion Operations
Review Emergency OBT Simulator Functionality
NanoRacks Cubesat Video Setup
Health Maintenance System (HMS) Vision Questionnaire
Multi-purpose Small Payload Rack (MSPR) /Group Combustion Module (GCM) Component Activation
CB/ISS CREW CONFERENCE
Meteor Shutter Close

Completed Task List Items
Carbon Dioxide Monitor (CDM) Data Download
Relocate CTB from Columbus to PMM
ESA Active Dosimeter Area Monitoring
Columbus System CD/DVD swap
FGB EDV Relocate
Dragon Cargo Transfer Operations
Carbon Dioxide Monitor (CDM) Set-up
Photo TV D4 Camera Sensor Cleaning
Transfer 66P Unpack

Ground Activities
All activities were completed unless otherwise noted.
NRCSD #10 Deploy
MESC Teardown support

Three-Day Look Ahead:
Tuesday, 03/07: Body Measures, PMA3 Relocate Preps, Auxin Transport, OBT EMER sim, OCT Eye Exam
Wednesday, 03/08: Eye Ultrasound, N3 Port CPA Install, NanoRack Module 9, APEX-4
Thursday, 03/09: CubeSate Deployer Removal, HXP Adapter Install, Fluid Shifts, EVA Tool Config, PMA3 Relocate Preps

QUICK ISS Status - Environmental Control Group:
Component - Status
Elektron - On
Vozdukh - Manual
[СКВ] 1 - SM Air Conditioner System ("SKV1") - On
[СКВ] 2 - SM Air Conditioner System ("SKV2") - Off
Carbon Dioxide Removal Assembly (CDRA) Lab - Standby
Carbon Dioxide Removal Assembly (CDRA) Node 3 - Operate
Major Constituent Analyzer (MCA) Lab - Operate
Major Constituent Analyzer (MCA) Node 3 - Operate
Oxygen Generation Assembly (OGA) - Process
Urine Processing Assembly (UPA) - Process
Trace Contaminant Control System (TCCS) Lab - Off
Trace Contaminant Control System (TCCS) Node 3 - Full Up

Please follow SpaceRef on Twitter and Like us on Facebook.